OpenStudy (goten77):

If i have y=(10x+5)/5 this can be re written as y= 2x+1 ????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=(10x+5 )/5 = 5(2x+1)/5=2x+1